OpenStudy (anonymous):

solve for x a. 1/e^x=2 b. ln(2x)=7

OpenStudy (anonymous):

\[\frac{1}{e^x}=2\iff e^x=\frac{1}{2}\] is a start

OpenStudy (anonymous):

and the times ln on both sides

OpenStudy (anonymous):

and then please

OpenStudy (anonymous):

satellite 73

OpenStudy (anonymous):

write in equivalent logarithmic form as \[x=\ln(\frac{1}{2})\]

OpenStudy (anonymous):

x=ln/2